1) What mass (in kg) of iron must be placed on top of the wood block so that the wood block is just barely submerged (the top of the block of wood is level with the surface of the water and the iron on top is in the air). The volume of the wood block is 0.00782 m^3. The density of the wood is 620 kg/m3 and the density of the water is 1000 kg/m3.

2) Most cars have hydraulic brakes. When a person pushes on the brake pedal, a force is applied to a small piston with a cross-sectional area of 1.4 cm2. By Pascal's Principle, the pressure is transmitted through the brake fluid to two larger pistons (each with an area of 9 cm2). The larger pistons push on the brake pads and the brake pads rub against the rotor (which stops the car). When a force of 10 N is applied by the brake pedal to the small piston, what force (in N) does each of the larger pistons apply to the rotor?

3) Consider a 2.85 m3 volume of argon gas at STP. STP means "Standard Temperature and Pressure", which are taken to be 0 ° C and 101,325 Pa, respectively. How many moles (in mol) of argon gas are there in this sample? How many atoms (no units) of argon gas are there in this sample?

4) A bridge is made of concrete slabs that are 32 m long at 27.6 ° C. What is the total length (in m) of each of the concrete slabs at 15.4 ° C? The Coefficient of Linear Expansion for concrete is 12 x 10-6 ° C-1. You must keep 5 SIGNIFICANT FIGURES in your answer.
Answer #1

1] Let mass of iron be m, Normal force on iron by wood = mg,

also normal force on wood by iron = mg downward

Now force equation : weight of wood + normal by iron - buoyant force = 0

rhowood*V*g + mg - rhowater*V*g = 0

620*0.00782 + m - 1000*0.00782 = 0

m = 1000*0.00782 - 620*0.00782

= 2.9716 kg answer